# SoK: The Design Paradigm of Safe and Secure Defaults

### v1

@@ Abstract
In security engineering, including software security engineering, there is a well-known design paradigm telling to prefer safe and secure defaults. The paper presents a systematization of knowledge (SoK) of this paradigm by the means of a systematic mapping study and a scoping review of relevant literature. According to the mapping and review, the paradigm has been extensively discussed, used, and developed further since the late 1990s. Partially driven by the insecurity of the Internet of things, the volume of publications has accelerated from the circa mid-2010s onward. The publications reviewed indicate that the paradigm has been adopted in numerous different contexts. It has also been expanded with security design principles not originally considered when the paradigm was initiated in the mid-1970s. Among the newer principles are an “off by default” principle, various overriding and fallback principles, as well as those related to the zero trust model. The review also indicates obstacles developers and others have faced with the paradigm.
@@ 1 Introduction
Table 1: The Eight Security Design Principles of Saltzer and Schroeder [134, quotations from pp. 1282–1283]
@@
Although the Saltzer’s and Schroeder’s paper is a modern classic in computer science, thus far, no systematic reviews have been conducted about its use in academic research. The review presented fills this knowledge gap. As will be seen, the design paradigm has been continuously discussed in the face of new technologies. It has also been extended and customized to meet requirements originating from new design problems. Last but by no means least, it has frequently been misused in practice.
@@
The focus on security design principles frames the paper toward security engineering. Alas, there are no universally agreed definitions for security engineering. In general, it can be seen to be about putting “security theory into security practice”, meaning that “a security engineer designs and makes systems that are protected against threats, i.e. forces to which systems may be subjected” [140, p. 59]. In the present context the notion about putting theory into practice is about adapting and applying the abstract design paradigm in more practical contexts. The framing toward security engineering is also important because it puts the paradigm into a context of engineering secure systems; hence, the research reviewed is also largely about defensive cyber security. Although there are papers about vulnerabilities in existing systems, whether hardware or software, the majority of papers reviewed are about building or improving systems, software, networks, protocols, or technologies in general with security and safety in mind.
@@ 2 Methodology
Even with the criteria, preliminary queries indicated a vast amount of peer reviewed literature. To deal with this severe feasibility obstacle, the final searches were restricted only to the ACM’s and IEEE’s electronic libraries. This restriction underlines the paper’s nature as a scoping review. Regarding the noted value contemplation before conducting a full systematic literature review, it can be tentatively concluded that not much additional value would be supposedly available because the $n=148$ reviewed papers (see Fig. 1) are already sufficient for conveying the relevant points raised and soon discussed. If further databases would be queried, it also remains unclear whether it would be possible to review the literature without restoring to natural language processing or bibliometrics.
@@
It can be also noted that the criteria were strictly followed in order to minimize subjectivity. Therefore, it is worth further remarking that there is a paper [19] that has likely used so-called tortured phrases [16] or something alike. To deal with such papers as well as the overall feasibility problem, it might be possible to evaluate the quality of papers, but the problem is that there is no universally agreed definition for paper quality [14]. This lack of a definition would presumably increase subjectivity. A potential solution would be to focus on national rankings of publication venues [40], journal impact factors, or related quantitative information, but also this choice would likely bias a literature sample because high-quality work is published also in low-prestige venues, and the other way around. These points notwithstanding, it can be remarked that the overall quality of the sample is good based on a subjective evaluation. This remark can be reinforced by taking a look at the publication venues; many, but not obviously all, papers were published in well-known or even top-ranked security conferences and journals.
Finally, the reviewing itself follows a thematic approach that is common in qualitative research [128, 129]. In other words, the goal is to capture and categorize the major themes present in the literature. The thematic approach is also necessary due to the large amount of literature. In practice, the reviewing was started by reading each paper’s abstract. The second step was about reading the portions in which defaults were discussed. If a more detailed reading was required to deduce about the defaults, a paper’s introduction and conclusions were first read. If these were not sufficient, as was the case with a few papers, a full reading was conducted, from the first to the last sentence.
@@ 3.2 Domains
Operating systems (including virtualization) [95, 60, 137, 167, 55, 149, 108, 136, 148, 85, 13, 161, 122, 73, 144, 172, 3, 168, 90]
@@
Although a thematic analysis is not well-suited for quantitative insights, it can be still concluded that embedded devices, including particularly the Internet of things (IoT) devices, have been the leading force behind the design paradigm—or, rather, its abuses. In fact, the bigram internet thing takes the fourth place in Fig. 3. This domain is also referred to as the Internet of vulnerable things in the literature [121]. Alternatively, the insecurity of “thingernets” is related to the vulnerabilities in the “thingabilities” and “thingertivity” of things [89]. To put the paper’s humor aside, the “thingernets” concept is illuminating in a sense that it has been estimated that it was around the early 2010s when more “thingabilities” than humans were connected to the Internet according to the non-sampled materials [35]. Thus, the IoT domain largely explains also the acceleration of papers discussing the design paradigm (cf. Fig. 2). The reason is the domain’s overall insecurity. In this regard, it is worth continuing the historical narrative by pointing out a couple of papers discussing the Mirai botnet built around compromised IoT devices [163, 77]. It was first discovered in 2016, and rang the alarm bells throughout the computing world.
@@ 3.3 Motivations
By and large, the whole paradigm has been motivated in the literature though a negation. In other words, as can be also concluded from Fig. 6, the starting point and motivation have usually been unsafe and insecure defaults. The motivations through a negation are expectedly particularly pronounced in the IoT domain. Some publications talk about an insecure default configuration problem, meaning that IoT devices are routinely shipped with extremely poor configurations with respect to best—or even elementary—cyber security practices [46]. As has been pointed out in numerous publications, these configuration mistakes include, but are not limited to, default or weak passwords. Regarding the other issues, the IoT domain is shaped by “extreme heterogeneity, mostly plug-and-play nature, computational limitations, improper patch management, unnecessary open ports, default or no security credentials, and extensive use of reusable open-source software” [146, p. 11224]. On one hand: as with the Mirai botnet, it is also worth emphasizing that the insecurity problems can escalate into larger, even society-wide problems. Although many of the publications analyzed, discussed, and elaborated the IoT domain with conventional consumer-grade devices, such as routers and modems, the same issues seem to plague even smart grid devices [2]. On the other hand: although IoT devices have likely exacerbated the problem, it is worth noting a publication from 2010 that discussed the same issues with default passwords [18]. Another publication from 2012 analyzed the insecure default settings in Windows and Linux operating systems at that time [135]. Thus, like with the design paradigm itself, the antonyms of unsafe and insecure defaults are nothing new as such in the computing world.
@@
Figure 5: New and Old Design Principles
Figure 6: Obstacles
The security issues in the IoT domain translate also into network protocols. In particular, numerous publications have analyzed and tried to improve the so-called MQTT protocol, which does not encrypt traffic by default. In this regard, the QUIC protocol has been seen as a promising alternative [90]. However, it should be noted that the issues with protocols go beyond the IoT and related domains; among other things, the same-origin policy for the world wide web has been seen to violate the design paradigm too [70]. In any case, cryptography is also a good example otherwise; among other things, some publications have analyzed insecure use of cryptographic libraries; hence, secure defaults should be provided in these. Insecure defaults are indirectly reflected also in observations that some cryptographic libraries are prone to side-channel attacks [12]. Regarding other common protocols: while the transport layer protocol (TLS) is the de facto one for the today’s world wide web, electronic mail is still unencrypted by default, which has motivated some to try to improve the situation [19]. Another example would be nowadays popular end-to-end encryption; people seem to lack knowledge about it and do not trust it [27]. As always with cryptography, a point has also been raised that encryption of IoT and other devices may cause obstacles for forensics and law enforcement investigations [170]. A further point is that the antonyms have not motivated only research on IoT, protocols, and cryptography. A good example would be the operating system domain; therein, some virtualization solutions have been seen to unnecessarily expose too many system calls [168]. Another example would be possibly insecure push notifications used in smartphones by default [109]. All these examples can be seen to belong to the domain of insecure and unsafe defaults, the antonyms of the design paradigm.
@@
Three publications used system monitoring, dynamic behavior of networks, and a risk of a single point of failure as motivations. Remotely conducted monitoring was seen to provide means to detect misconfigured security settings [167], traffic locality and dynamics exhibited were argued to contradict often coarse-grained and static default network configurations [65], and a dynamic conflict arbiter was designed to prevent single points of failure [25]. It can be noted that in the non-sampled literature single points of failure are related to the classical security design principle of defense in depth [153]. There was also an odd category labeled as opt-in. For instance, an argument was raised against the C language in that its use “rely too much on disciplined opt-in; there is still no safe default” [76, p. 243]. The category is closely related to the non-technical motivations labeled as human factors in Fig. 6.
@@
Among other things, they may lack knowledge, they may defer changes to settings due to cognitive biases, including a bias that a default setting conveys information about what is reasonable, or they may lack technical skills to make changes [142]. There is also a so-called status quo bias; people allegedly fear that changing security configurations break existing functionality [142, 147]. Though, the incentives or biases may also be rational in a sense; people allegedly stick to default security features in operating systems because they are provided free of charge [108]. Whatever the actual reasons might be, some [30] or most [8] people never modify default configurations. Passwords are a good example. For instance, one study found that “an outstanding 83.2% of the students have not changed their passwords” [11, p. 1431]. However, similar claims do not affect only end-users. Also software developers tend to use default security features offered by a given provider [102]. Furthermore, there is a phishing-related angle to this reasoning; even users who follow good security practices may opt for insecure solutions when presented a choice during their work flows [171]. There are a couple of problems in these claims and the associated reasoning about human factors related to default settings.
@@ 3.4 Design Principles
The literature has presented numerous security design principles, some of which are relatively straightforward security design principles and some others more general ideas about good security practices to adopt. All of these are on the technical side; none of the papers sampled provided feasible solutions on how the problems with human factors could be resolved, provided that generic points raised about improved education and security awareness are not counted. The principles are summarized in Fig. 6. The figure can be unpacked and elaborated with ten points.
@@
First, a few papers have operated with high-level design principles. These include the Saltzer’s and Schroeder’s economy principle. For instance, “the recent emphasis on simplifying APIs (and choosing secure defaults) has provided improvement; we endorse continuing in this direction” [1, p. 168]. The economy principle has also been correlated with the psychological acceptability principle in the literature; “designs should be as simple as possible and provide users with a secure default path-of-least-resistance to complete authentication” [103, p. 203]. The notion about path-of-least-resistance is also known as a user-buy-in security design principle in the non-sampled literature [153]. In addition to the economy principle, a general abstraction design principle and a principle of isolation can be seen to belong to the category of high-level design principles. In terms of the former, abstracting security features has been argued to reduce the likelihood that developers will misuse the features, possibly introducing vulnerabilities along the way [81]. Then, isolation has been used to restrict computer communication between entities [52], and more generally, “an isolation-by-default policy results in a significant increase in user security” [32, p. 1342]. It should be remarked that isolation as a default is again nothing new as such; in the general (non-sampled) security literature it is essentially also behind the compartmentalization design principle [153]. Regardless of a terminology, in the literature reviewed isolation has gained newfound interest due to the emergence of virtualization, cloud computing, and associated technologies and services.
@@
Second, as already noted, quite a few publications have operated with access controls and the least privilege principle considered already by Saltzer and Schroeder. The works in this domain often use apt slogans to motivate their designs. “Deny by default” [20, p. 509] is a good example. Although the basic idea remains the same, the slogans vary from a context to another. For instance, publications dealing with input and output operations have justified their designs by security policy notions such as a “fail-safe default of no write access” [80, p. 119] or a default “to automatically deny the read-write” [7, p. 260]. These access control slogans and security policy dictates align with the principle of blacklists and whitelists. Before continuing, it should be noted that the former violates the Saltzer’s and Schroeder’s original fail-safe default principle, whereas the latter conforms with it. Among other things, whitelists have been used to restrict parameters supplied to cryptographic libraries [75], whereas “billions of devices benefit from the blacklist enabled by default” [31, p. 4345] in browsers. This quotation demonstrates that it is not always possible to fulfill the ideal of denying by default; there are billions (or more) legitimate websites, and thus no one can curate a whitelist in practice.
@@
Third, a newer principle of disabling functionality is prominent in the literature. The explanation is partially related to the overall insecurity in the IoT domain. A few illuminating quotations help to again elaborate this new principle. “Making communication default-off has tangible security benefits” [61, p. 117]. This “off-by-default approach” [43, p. 197] is commonly endorsed in the computer network research domain; all communication is denied by default [39, p. 321] and computer networks have been designed so that they “only admit the expected traffic, by default treating the rest as unwanted traffic” [74, p. 1]. In fact, recent “proposals for capabilities-based networks have provided some ideas on the fundamental shift in the design philosophy of networks by moving from the Internets ‘on-by-default’ principle to an ‘off-by-default’ assumption” [162, p. 1931]. However, the principle generalizes to other domains as well. For instance, web security has been argued to improve from “disabling the behavior by default in web browsers” [69, p. 673], including with respect to browser plugins [72]. Furthermore, the off-by-default dictum has been adopted as a general argument about delivering software products; they “should really be distributed with access disabled until administrators explicitly customize the access control policy and mechanisms for their organization” [28, p. 10]. Finally, the new principle also aligns in the literature with the economy principle; a design “aims for true minimality in the sense that nothing should be included by default that the service does not explicitly need” [13, p. 251]. It is not difficult to come up with other slogans aligning with the principle—it could be described also with phrasing such as “disable by default, then enable”, “close by default, then open”, and so forth and so on.
@@
Fourth, the off-by-default dictum should not be generalized to everything; the literature also emphasizes a principle of turning security features on by default. In addition to technical designs and their implementations, including those mandating an execution of specific instructions [166] and those related to enabling sanitization routines in web applications [115], the principle aligns with the human factor side. The incentives discussed earlier are visible in observations that study participants “were happy to use two-factor authentication as a secure default that is set up at registration, while they would not go through the steps to set it up afterwards” [54, p. 12]. Thus, also this principle aligns with the psychological acceptability principle; people should be encouraged to adopt good security practices, whether via education or user interface designs.
@@
Fifth, there is a principle related to overriding functionality enabled by default. For instance, a design hooked calls to existing APIs in order to enforce validation of host names and certificates [9]. Another good example involves computer networking; a study designed a system via which “the default credential is exchanged for a credential that is suitable for the service in question” [155, p. 150]. Though, an argument can be also raised that overriding things should not be necessary in the ideal world because it indicates that something is non-optimal in a design that is being overridden.
@@
Sixth, to some extent, the already noted clean-up design principle aligns with a principle to prevent leaks of different kinds. Such information or data leaks vary from a context to another. A good example would be enforcing “best practices when enabling logs, ensuring the server administrator really intended to expose this information” [96, p. 4]. Another example would be plugging leaks in honeypots in order to improve deception and prevent detection by adversaries [17]. In the general, non-sampled literature the detection techniques are known as anti-honeypot and anti-introspection methods [151]. As could be expected, the leak prevention principle has also been adopted to protect the privacy of users by default [165, 107]. Thus, leaks are a good example on how security design principles sometimes extend to other domains as well.
@@
Seventh, there are design principles related to automation and generating defaults. Regarding the former, examples include automation of network security configurations [120], providing personalized security settings [143], automation of virtual machine configurations [148], and automatic provision of certificates to end-users [104]. Regarding the default generating principle, a paper presented a technique to ensure a secure default option by generating strong passwords automatically for users [56]. While a full automation of security configurations may be challenging and thus also risky, the latter idea would seem sensible for also fixing some of the basic issues with IoT devices; user names and strong passwords could be generated at a factory and perhaps printed to stickers placed at the bottoms of all IoT devices manufactured and shipped.
@@
Eight, there is a design principle related to explicitly guiding developers about security. For instance, a paper designed a fail-safe default that forces developers to check and validate their configurations [111]. Another paper dealing with databases designed a solution “forcing security modellers to think about any cycles in a schema, and add explicit constraints to weaken the schema only where they believe this is safe” [150, p. 113]. This principle aligns with the earlier points about secure defaults for cryptographic libraries, further strengthening the defaults with explicit constraints.
@@
Ninth, there are design principles about robust fallback and termination. Some of these principles are on the safety side and the domain of cyber-physical systems. For instance, a custom solution for unmanned aerial vehicles (UAVs) was compared against a “default mitigation strategy” involving opening a “parachute when a failure is detected” [50]. Another paper operating in the UAV domain discussed fail-safe defaults of either landing or returning to home [21]. That said, sometimes there is no other option but to gracefully terminate an execution; “stopping locomotion when an error is detected is a safe default action” [48, p. 150] in some robotics applications. Closer to the domain of software and systems, the fallback design principle often involves reverting to a last known good or valid state in case of a malfunction [164]. Some of the fallback designs also correlate with the other principles outlined. Among these is isolation; “monitor the system and rapidly adapt when conditions change, falling back to strict isolation as the safe default” [83, p. 351]. Analogously, a “system responded by disabling all network ports and defaulted to the safe-operation mode established previously” [156, p. 1]. While these and other fallback and termination designs were not considered by Saltzer and Schroeder, they too belong to the principle of fail-safe defaults.

### v2

@@ Abstract
In security engineering, including software security engineering, there is a well-known design paradigm telling to prefer safe and secure defaults. The paper presents a systematization of knowledge (SoK) of this paradigm by the means of a systematic mapping study and a scoping review of relevant literature. According to the mapping and review, the paradigm has been extensively discussed, used, and developed further since the late 1990s. Partially driven by the insecurity of the Internet of things, the volume of publications has accelerated from the circa mid-2010s onward. The publications reviewed indicate that the paradigm has been adopted in numerous different contexts. It has also been expanded with security design principles not originally considered when the paradigm was initiated in the mid-1970s. Among the newer principles are an “off by default” principle, various overriding and fallback principles, as well as those related to the zero trust model. The review also indicates problems developers and others have faced with the paradigm.
@@ 1 Introduction
Table 1: The Eight Security Design Principles of Saltzer and Schroeder [1, quotations from pp. 1282–1283]
@@
The Saltzer’s and Schroeder’s paper is a modern classic in computer science. Yet, thus far, no systematic reviews have been conducted about its use in academic research, although there are some existing reviews that come close [2]. The review presented fills the apparent knowledge gap. In addition, the review fulfills other common functions of literature reviews, including the identification of trends and patterns, clarification of complex concepts, and support for education through cataloging important or even foundational knowledge [3]. Regarding trends, as will be seen, the design paradigm has been continuously discussed in the face of new technologies. It has also been extended and customized to meet requirements originating from new design problems. Furthermore, it has frequently been misused in practice. This misuse raises also the review’s practical relevance. As will be pointed out later on in the concluding Section 4, secure defaults are mandated by recent cyber security regulations. Thus, the SoK is beneficial also for practitioners and regulators who both may find new insights from the knowledge cataloged.
@@
The focus on security design principles frames the paper toward security engineering. Alas, there is no single universally agreed definition for security engineering. In general, it can be seen to be about putting “security theory into security practice”, meaning that “a security engineer designs and makes systems that are protected against threats, i.e. forces to which systems may be subjected” [6, p. 59]. In the present context the notion about putting theory into practice is about adapting and applying the abstract design paradigm in more practical contexts. The framing toward security engineering is also important because it puts the paradigm into a context of engineering secure systems; hence, the research reviewed is also largely about defensive cyber security. Although there are papers about vulnerabilities in existing systems, whether hardware or software, the majority of papers reviewed are about building or improving systems, software, networks, protocols, or technologies in general with security and safety in mind.
@@ 2 Methodology
Even with the criteria, preliminary queries indicated a vast amount of peer reviewed literature. To deal with this severe feasibility obstacle, the final searches were restricted only to the ACM’s and IEEE’s electronic libraries. This restriction underlines the paper’s nature as a scoping review. Regarding the noted value contemplation before conducting a full systematic literature review, it can be tentatively concluded that not much additional value would be supposedly available because the $n=148$ reviewed papers (see Fig. 1) are already sufficient for conveying the relevant points raised and soon discussed. If further databases would be queried, it also remains unclear whether it would be possible to review the literature without resorting to natural language processing or bibliometrics.
@@
It can be also noted that the criteria were strictly followed in order to minimize subjectivity. Therefore, it is worth further remarking that there is a paper [10] that has likely used so-called tortured phrases [11] or something alike. To deal with such papers as well as the overall feasibility problem, it might be possible to evaluate the quality of papers [12], but the problem is that there is no universally agreed definition for paper quality [13]. This lack of a definition would presumably increase subjectivity. A potential solution would be to focus on national rankings of publication venues [14], journal impact factors, or related quantitative information, but also this choice would likely bias a literature sample because high-quality work is published also in low-prestige venues, and the other way around. These points notwithstanding, it can be remarked that the overall quality of the sample is good based on a subjective evaluation. This remark can be reinforced by taking a look at the publication venues; many, but not obviously all, papers were published in well-known or even top-ranked security conferences and journals.
Finally, the reviewing itself follows a thematic approach that is common in qualitative research [15, 16]. In other words, the goal is to capture and categorize the major themes present in the literature. The thematic approach is also necessary due to the large amount of literature sampled. The themes were constructed by first identifying a large number of distinct categories, which were then narrowed and sharpened by collating and merging of overlapping, overcasting, or otherwise fuzzy categories  [16, 17]. In practice, the identification of themes and the reviewing in general were started by reading each paper’s abstract. The second step was about reading the portions in which defaults were discussed. If a more detailed reading was required to deduce about the defaults, a paper’s introduction and conclusions were first read. If these were not sufficient, as was the case with a few papers, a full reading was conducted, from the first to the last sentence.
@@ 3 Review
In what follows, the review is presented in five subsections. The first subsection presents a few quantitative observations about the literature sample. The two subsequent subsections elaborate the computing domains in which the papers have operated together with their motivations. The design principles behind the paradigm are discussed in the next subsection. The final subsection briefly discusses some problems identified in the literature.
@@ 3.2 Domains
Operating systems (including virtualization) [105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122]
@@
Although a thematic analysis is not well-suited for quantitative insights, it can be still concluded that embedded devices, including particularly the Internet of things (IoT) devices, have been the leading force behind the design paradigm—or, rather, its abuses. In fact, the bigram internet thing takes the fifth place in Fig. 3. This domain is also referred to as the Internet of vulnerable things in the literature [54]. Alternatively, the insecurity of “thingernets” is related to the vulnerabilities in the “thingabilities” and “thingertivity” of things [72]. To put the paper’s humor aside, the “thingernets” concept is illuminating in a sense that it has been estimated that it was around the early 2010s when more “thingabilities” than humans were connected to the Internet according to the non-sampled materials [166]. Thus, the IoT domain largely explains also the acceleration of papers discussing the design paradigm (cf. Fig. 2). The reason is the domain’s overall insecurity. In this regard, it is worth continuing the historical narrative by pointing out a couple of papers discussing the Mirai botnet built around compromised IoT devices [70, 74]. It was first discovered in 2016, and rang the alarm bells throughout the computing world.
@@ 3.3 Motivations
By and large, the whole paradigm has been motivated in the literature though a negation. In other words, as can be also concluded from Fig. 6, which visualizes the motivating themes used in the papers, the starting point and motivation have usually been unsafe and insecure defaults. The motivations through a negation are expectedly particularly pronounced in the IoT domain. Some publications talk about an insecure default configuration problem, meaning that IoT devices are routinely shipped with extremely poor configurations with respect to best—or even elementary—cyber security practices [49]. As has been pointed out in numerous publications, these configuration mistakes include, but are not limited to, default or weak passwords. Regarding the other issues, the IoT domain is shaped by “extreme heterogeneity, mostly plug-and-play nature, computational limitations, improper patch management, unnecessary open ports, default or no security credentials, and extensive use of reusable open-source software” [73, p. 11224]. On one hand: as with the Mirai botnet, it is also worth emphasizing that the insecurity problems can escalate into larger, even society-wide problems. Although many of the publications analyzed, discussed, and elaborated the IoT domain with conventional consumer-grade devices, such as routers and modems, the same issues seem to plague even smart grid devices [71]. On the other hand: although IoT devices have likely exacerbated the problem, it is worth noting a publication from 2010 that discussed the same issues with default passwords [65]. Another publication from 2012 analyzed the insecure default settings in Windows and Linux operating systems at that time [103]. Thus, like with the design paradigm itself, the antonyms of unsafe and insecure defaults are nothing new as such in the computing world.
@@
Figure 5: Design Principles
Figure 6: Problems
The security issues in the IoT domain translate also into network protocols. In particular, numerous publications have analyzed and tried to improve the so-called MQTT protocol, which does not encrypt traffic by default. In this regard, the QUIC protocol has been seen as a promising alternative [122]. However, it should be noted that the issues with protocols go beyond the IoT and related domains; among other things, the same-origin policy for the world wide web has been seen to violate the design paradigm too [161]. In any case, cryptography is also a good example otherwise; among other things, some publications have analyzed insecure use of cryptographic libraries; hence, secure defaults should be provided in these.
Insecure defaults are indirectly reflected also in observations that some cryptographic libraries are prone to side-channel attacks [38]. Regarding other common protocols: while the transport layer protocol (TLS) is the de facto one for the today’s world wide web, electronic mail is still unencrypted by default, which has motivated some to try to improve the situation [10]. Another example would be nowadays popular end-to-end encryption; people seem to lack knowledge about it and do not trust it [39]. As always with cryptography, a point has also been raised that encryption of IoT and other devices may cause obstacles for forensics and law enforcement investigations [69]. A further point is that the antonyms have not motivated only research on IoT, protocols, and cryptography. A good example would be the operating system domain; therein, some virtualization solutions have been seen to unnecessarily expose too many system calls [121]. Another example would be possibly insecure push notifications used in smartphones by default [146]. All these examples can be seen to belong to the domain of insecure and unsafe defaults, the antonyms of the design paradigm.
@@
Three publications used system monitoring, dynamic behavior of networks, and a risk of a single point of failure as motivations. Remotely conducted monitoring was seen to provide means to detect misconfigured security settings [108], traffic locality and dynamics exhibited were argued to contradict often coarse-grained and static default network configurations [86], and a dynamic conflict arbiter was designed to prevent single points of failure [30]. It can be noted that in the non-sampled literature single points of failure are related to the classical security design principle of defense in depth [21]. There was also an odd category labeled as opt-in. For instance, an argument was raised against the C language in that its use “rely too much on disciplined opt-in; there is still no safe default” [127, p. 243]. The category is closely related to the non-technical motivations labeled as human factors in Fig. 6.
@@
Among other things, they may lack knowledge, they may defer changes to settings due to cognitive biases, including a bias that a default setting conveys information about what is reasonable, or they may lack technical skills to make changes [76]. There is also a so-called status quo bias; people allegedly fear that changing security configurations break existing functionality [76, 23]. Though, the incentives or biases may also be rational in a sense; people allegedly stick to default security features in operating systems because they are provided free of charge [111]. Whatever the actual reasons might be, some [139] or most [147] people never modify default configurations. Passwords are a good example. For instance, one study found that “an outstanding 83.2% of the students have not changed their passwords” [124, p. 1431]. However, similar claims do not affect only end-users. Also software developers tend to use default security features offered by a given provider [145]. Furthermore, there is a phishing-related angle to this reasoning; even users who follow good security practices may opt for insecure solutions when presented a choice during their work flows [138]. There are a couple of problems in these claims and the associated reasoning about human factors related to default settings.
@@ 3.4 Design Principles
The literature has presented numerous security design principles, some of which are relatively straightforward security design principles and some others more general ideas about good security practices to adopt. All of these are on the technical side; none of the papers sampled provided feasible and plausible ideas on how the problems with human factors could be resolved, provided that generic points raised about improved education and security awareness are not counted. On these notes, the design principles are summarized in Fig. 6.
@@
A few papers have operated with high-level design principles. These include the Saltzer’s and Schroeder’s economy principle. For instance, “the recent emphasis on simplifying APIs (and choosing secure defaults) has provided improvement; we endorse continuing in this direction” [37, p. 168]. The economy principle has also been correlated with the psychological acceptability principle in the literature; “designs should be as simple as possible and provide users with a secure default path-of-least-resistance to complete authentication” [154, p. 203]. The notion about path-of-least-resistance is also known as a user-buy-in security design principle in the non-sampled literature [21]. In addition to the economy principle, a general abstraction design principle and a principle of isolation can be seen to belong to the category of high-level design principles. In terms of the former, abstracting security features has been argued to reduce a likelihood that developers will misuse the features, possibly introducing vulnerabilities along the way [151]. Then, isolation has been used to restrict computer communication between entities [149], and more generally, “an isolation-by-default policy results in a significant increase in user security” [47, p. 1342]. It should be remarked that isolation as a default is again nothing new as such; in the non-sampled security literature it is essentially also behind the compartmentalization design principle [21]. Regardless of a terminology, in the literature reviewed isolation has gained a newfound interest due to the emergence of virtualization, cloud computing, and associated technologies and services.
@@
As already noted, quite a few publications have operated with access controls and the least privilege principle considered already by Saltzer and Schroeder. The works in this domain often use apt slogans to motivate their designs. “Deny by default” [45, p. 509] is a good example. Although the basic idea remains the same, the slogans vary from a context to another. For instance, publications dealing with input and output operations have justified their designs by security policy notions such as a “fail-safe default of no write access” [44, p. 119] or a default “to automatically deny the read-write” [46, p. 260]. These access control slogans and security policy dictates align with a principle of blacklists and whitelists. Before continuing, it should be noted that the former violates the Saltzer’s and Schroeder’s original fail-safe default principle, whereas the latter conforms with it. Among other things, whitelists have been used to restrict parameters supplied to cryptographic libraries [32], whereas “billions of devices benefit from the blacklist enabled by default” [155, p. 4345] in web browsers. This quotation demonstrates that it is not always possible to fulfill the ideal of denying by default; there are billions (or more) of legitimate websites, and thus no one can curate a whitelist in practice.
@@
A principle of disabling functionality is prominent in the literature. The explanation is partially related to the overall insecurity in the IoT domain. A few illuminating quotations help to again elaborate this principle. “Making communication default-off has tangible security benefits” [50, p. 117]. This “off-by-default approach” [85, p. 197] is commonly endorsed in the computer network research domain; all communication is denied by default [90, p. 321] and computer networks have been designed so that they “only admit the expected traffic, by default treating the rest as unwanted traffic” [101, p. 1]. In fact, recent “proposals for capabilities-based networks have provided some ideas on the fundamental shift in the design philosophy of networks by moving from the Internets ‘on-by-default’ principle to an ‘off-by-default’ assumption” [91, p. 1931]. However, the principle generalizes to other domains as well. For instance, web security has been argued to improve from “disabling the behavior by default in web browsers” [157, p. 673], including with respect to browser plugins [158]. Furthermore, the off-by-default dictum has been adopted as a general argument about delivering software products; they “should really be distributed with access disabled until administrators explicitly customize the access control policy and mechanisms for their organization” [152, p. 10]. Finally, the principle also aligns in the literature with the economy principle; a design “aims for true minimality in the sense that nothing should be included by default that the service does not explicitly need” [115, p. 251]. It is not difficult to come up with other slogans aligning with the principle—it could be described also with phrasing such as “disable by default, then enable”, “close by default, then open”, and so forth and so on.
@@
The off-by-default dictum should not be generalized to everything; the literature also emphasizes a principle of turning security features on by default. In addition to technical designs and their implementations, including those mandating an execution of specific instructions [22] and those related to enabling sanitization routines in web applications [153], the principle aligns with the human factor side. The incentives discussed earlier are visible in observations that study participants “were happy to use two-factor authentication as a secure default that is set up at registration, while they would not go through the steps to set it up afterwards” [143, p. 12]. Thus, also this principle aligns with the psychological acceptability principle; people should be encouraged to adopt good security practices, whether via education or user interface designs.
@@
There is also a principle related to overriding functionality enabled by default. For instance, a design hooked calls to existing APIs in order to enforce validation of host names and certificates [35]. Another good example involves computer networking; a study designed a system via which “the default credential is exchanged for a credential that is suitable for the service in question” [162, p. 150]. Though, an argument can also be raised that overriding things should not be necessary in the ideal world because it indicates that something is non-optimal in a design that is being overridden.
@@
To some extent, the already noted clean-up design principle aligns with a principle to prevent leaks of different kinds. Such information or data leaks vary from a context to another. A good example would be enforcing “best practices when enabling logs, ensuring the server administrator really intended to expose this information” [83, p. 4]. Another example would be plugging leaks in honeypots in order to improve deception and prevent detection by adversaries [81]. In the general, non-sampled literature the detection techniques are known as anti-honeypot and anti-introspection methods [167]. As could be expected, the leak prevention principle has also been adopted to protect the privacy of users by default [156, 140]. Thus, leaks are a good example on how security design principles sometimes extend to other domains as well.
@@
Then, there are design principles related to automation and generating defaults. Regarding the former, examples include automation of network security configurations [96], providing personalized security settings [24], automation of virtual machine configurations [113], and automatic provision of certificates to end-users [40]. Regarding the default generating principle, a paper presented a technique to ensure a secure default option by generating strong passwords automatically for users [123]. While a full automation of security configurations may be challenging and thus also risky, the latter idea would seem sensible for also fixing some of the basic issues with IoT devices; user names and strong passwords could be generated at a factory and perhaps printed to stickers placed at the bottoms of all IoT devices manufactured and shipped.
@@
There is a further design principle related to explicitly guiding developers about security. For instance, a paper designed a fail-safe default that forces developers to check and validate their configurations [33]. Another paper dealing with databases designed a solution “forcing security modellers to think about any cycles in a schema, and add explicit constraints to weaken the schema only where they believe this is safe” [41, p. 113]. This principle aligns with the earlier points about secure defaults for cryptographic libraries, further strengthening the defaults with explicit constraints.
@@
Furthermore, there are design principles about robust fallback and termination. Some of these principles are on the safety side and the domain of cyber-physical systems. For instance, a custom solution for unmanned aerial vehicles (UAVs) was compared against a “default mitigation strategy” involving opening a “parachute when a failure is detected” [28]. Another paper operating in the UAV domain discussed a fail-safe default of either landing or returning to home [29]. That said, sometimes there is no other option but to gracefully terminate an execution; “stopping locomotion when an error is detected is a safe default action” [25, p. 150] in some robotics applications. Closer to the domain of software and systems, the fallback design principle often involves reverting to a last known good or valid state in case of a malfunction [84]. Some of the fallback designs also correlate with the other principles outlined. Among these is isolation; “monitor the system and rapidly adapt when conditions change, falling back to strict isolation as the safe default” [82, p. 351]. Analogously, a “system responded by disabling all network ports and defaulted to the safe-operation mode established previously” [26, p. 1]. While these and other fallback and termination designs were not considered by Saltzer and Schroeder, they too can be seen to belong to the principle of fail-safe defaults.